### v2

# MM-LLMs: Recent Advances in MultiModal Large Language Models

Abstract
In the past year, MultiModal Large Language Models (MM-LLMs) have undergone substantial advancements, augmenting off-the-shelf LLMs to support MM inputs or outputs via cost-effective training strategies. The resulting models not only preserve the inherent reasoning and decision-making capabilities of LLMs but also empower a diverse range of MM tasks. In this paper, we provide a comprehensive survey aimed at facilitating further research of MM-LLMs. Specifically, we first outline general design formulations for model architecture and training pipeline. Subsequently, we provide brief introductions of existing MM-LLMs, each characterized by its specific formulations. Additionally, we review the performance of MM-LLMs on mainstream benchmarks and summarize key training recipes to enhance the potency of MM-LLMs. Lastly, we explore promising directions for MM-LLMs while concurrently maintaining a real-time tracking website111https://mm-llms.github.io for the latest developments in the field. We hope that this survey contributes to the ongoing advancement of the MM-LLMs domain.
MM-LLMs: Recent Advances in MultiModal Large Language Models
Duzhen Zhang1111Equal contributions. , Yahan Yu2111Equal contributions. , Chenxing Li1 , Jiahua Dong3222Corresponding authors. , Dan Su1, Chenhui Chu2222Corresponding authors. and Dong Yu1 1Tencent AI Lab 2Kyoto University 3Shenyang Institute of Automation, Chinese Academy of Sciences scoutzhang@tencent.com, yahan@nlp.ist.i.kyoto-u.ac.jp
1 Introduction
MultiModal (MM) pre-training research has witnessed significant advancements in recent years, consistently pushing the performance boundaries across a spectrum of downstream tasks Li et al. (2020); Akbari et al. (2021); Fang et al. (2021); Yan et al. (2021); Li et al. (2021); Radford et al. (2021); Li et al. (2022); Zellers et al. (2022); Zeng et al. (2022b); Yang et al. (2022); Wang et al. (2022a, b). However, as the scale of models and datasets continues to expand, traditional MM models incur substantial computational costs, particularly when trained from scratch. Recognizing that MM research operates at the intersection of various modalities, a logical approach is to capitalize on readily available pre-trained unimodal foundation models, with a special emphasis on powerful Large Language Models (LLMs) OpenAI (2022). This strategy aims to mitigate computational expenses and enhance the efficacy of MM pre-training, leading to the emergence of a novel field: MM-LLMs.
MM-LLMs harness LLMs as the cognitive powerhouse to empower various MM tasks. LLMs contribute desirable properties like robust language generation, zero-shot transfer capabilities, and In-Context Learning (ICL). Concurrently, foundation models in other modalities provide high-quality representations. Considering foundation models from different modalities are individually pre-trained, the core challenge facing MM-LLMs is how to effectively connect the LLM with models in other modalities to enable collaborative inference. The predominant focus within this field has been on refining alignment between modalities and aligning with human intent via a MM Pre-Training (PT) + MM Instruction-Tuning (IT) pipeline.
With the debut of GPT-4(Vision) OpenAI (2023) and Gemini Team et al. (2023), showcasing impressive MM understanding and generation capabilities, a research fervor on MM-LLMs has been sparked. Initial research primarily focuses on MM content comprehension and text generation like (Open)Flamingo Alayrac et al. (2022); Awadalla et al. (2023), BLIP-2 Li et al. (2023c), Kosmos-1 Huang et al. (2023c), LLaVA/LLaVA-1.5 Liu et al. (2023e, d), MiniGPT-4 Zhu et al. (2023a), MultiModal-GPT Gong et al. (2023), VideoChat Li et al. (2023d), Video-LLaMA Zhang et al. (2023e), IDEFICS IDEFICS (2023), Fuyu-8B Bavishi et al. (2023), and Qwen-Audio Chu et al. (2023b). In pursuit of MM-LLMs capable of both MM input and output Aiello et al. (2023), some studies additionally explore the generation of specific modalities, such as Kosmos-2 Peng et al. (2023) and MiniGPT-5 Zheng et al. (2023b) introducing image generation, and SpeechGPT Zhang et al. (2023a) introducing speech generation. Recent research endeavors have focused on mimicking human-like any-to-any modality conversion, shedding light on the path to artificial general intelligence. Some efforts aim to amalgamate LLMs with external tools to reach an approaching ‘any-to-any’ MM comprehension and generation, such as Visual-ChatGPT Wu et al. (2023a), ViperGPT Surís et al. (2023), MM-REACT Yang et al. (2023), HuggingGPT Shen et al. (2023), and AudioGPT Huang et al. (2023b). Conversely, to mitigate propagated errors in the cascade system, initiatives like NExT-GPT Wu et al. (2023d) and CoDi-2 Tang et al. (2023b) have developed end-to-end MM-LLMs of arbitrary modalities. The timeline of MM-LLMs is depicted in Figure 1.
In this paper, we present a comprehensive survey aimed at facilitating further research of MM-LLMs. To provide readers with a holistic understanding of MM-LLMs, we initially delineate general design formulations from model architecture (Section 2) and training pipeline (Section 3). We break down the general model architecture into five components: Modality Encoder (Section 2.1), Input Projector (Section 2.2), LLM Backbone (Section 2.3), Output Projector (Section 2.4), and Modality Generator (Section 2.5). The training pipeline elucidates how to enhance a pre-trained text-only LLM to support MM input or output, primarily consisting of two stages: MM PT (Section 3.1) and MM IT (Section 3.2). In this section, we also provide a summary of mainstream datasets for MM PT and MM IT. Next, we engage in discussions of State-of-the-Art (SOTA) MM-LLMs, each characterized by specific formulations, and summarize their development trends in Section 4. In Section 5, we comprehensively review the performance of major MM-LLMs on mainstream benchmarks and distill key training recipes to enhance the efficacy of MM-LLMs. In Section 6, we offer promising directions for MM-LLMs research. Moreover, we have established a website (https://mm-llms.github.io) to track the latest progress of MM-LLMs and facilitate crowd-sourcing updates. Finally, we summarize the entire paper in Section 7 and discuss related surveys on MM-LLMs in Appendix A. We aspire for our survey to aid researchers in gaining a deeper understanding of this field and to inspire the design of more effective MM-LLMs.
2 Model Architecture
In this section, we provide a detailed overview of the five components comprising the general model architecture, along with the implementation choices for each component, as illustrated in Figure 2. MM-LLMs that emphasize MM understanding only include the first three components. During training, Modality Encoder, LLM Backbone, and Modality Generator are generally maintained in a frozen state. The primary optimization emphasis is on Input and Output Projectors. Given that Projectors are lightweight components, the proportion of trainable parameters in MM-LLMs is notably small compared to the total parameter count (typically around %). The overall parameter count is contingent on the scale of the core LLM utilized in the MM-LLMs. As a result, MM-LLMs can be efficiently trained to empower various MM tasks.
2.1 Modality Encoder
The Modality Encoder (ME) is tasked with encoding inputs from diverse modalities to obtain corresponding features , formulated as follows:
Various pre-trained encoder options exist for handling different modalities, where can be image, video, audio, 3D, or etc. Next, we will offer a concise introduction organized by modality.
Visual Modality
For images, there are generally four optional encoders: NFNet-F6 Brock et al. (2021), ViT Dosovitskiy et al. (2020), CLIP ViT Radford et al. (2021), and Eva-CLIP ViT Fang et al. (2023). NFNet-F6 is a normalizer-free ResNet He et al. (2016), showcasing an adaptive gradient clipping technique that allows training on extensively augmented datasets while achieving SOTA levels of image recognition. ViT applies the Transformer Vaswani et al. (2017) to images by first dividing the image into patches. It then undergoes linear projection to flatten the patches, followed by encoding via multiple Transformer blocks. CLIP ViT builds connections between text and images, comprising a ViT and a text encoder. Utilizing a vast amount of text-image pairs, it optimizes ViT by contrastive learning, treating paired text and images as positive samples and others as negative ones. Its Eva version stabilizes the training and optimization process of the massive CLIP, offering new directions in expanding and accelerating the expensive training of MM base models. For videos, they can be uniformly sampled to frames, undergoing the same pre-processing as images.
Audio Modality
is typically encoded by C-Former Chen et al. (2023b), HuBERT Hsu et al. (2021), BEATs Chen et al. (2023f), and Whisper Radford et al. (2023). C-Former employs the CIF alignment mechanism Dong and Xu (2020); Zhang et al. (2022a) for sequence transduction and a Transformer to extract audio features. HuBERT is a self-supervised speech representation learning framework based on BERT Kenton and Toutanova (2019), achieved by the masked prediction of discrete hidden units. BEATs is an iterative audio pre-training framework designed to learn Bidirectional Encoder representations from Audio Transformers.
3D Point Cloud Modality
is typically encoded by ULIP-2 Salesforce (2022); Xu et al. (2023a, b) with a PointBERT Yu et al. (2022) backbone.
Moreover, to handle numerous heterogeneous modal encoders, some MM-LLMs, particularly any-to-any ones, use ImageBind Girdhar et al. (2023), a unified encoder covering six modalities, including image, video, text, audio, heat map, etc.
2.2 Input Projector
The Input Projector is tasked with aligning the encoded features of other modalities with the text feature space . The aligned features as prompts are then fed into the LLM Backbone alongside the textual features . Given -text dataset , the goal is to minimize the -conditioned text generation loss :
where .
The Input Projector can be achieved directly by a Linear Projector or Multi-Layer Perceptron (MLP), i.e., several linear projectors interleaved with non-linear activation functions. There are also more complex implementations like Cross-attention, Q-Former Li et al. (2023c), or P-Former Jian et al. (2023). Cross-attention uses a set of trainable vectors as queries and the encoded features as keys to compress the feature sequence to a fixed length. The compressed representation is then fed directly into the LLM Bai et al. (2023b) or further used for X-text cross-attention fusion Alayrac et al. (2022). Q-Former extracts relevant features from , and the selected features are then used as prompts . Meanwhile, P-Former generates ‘reference prompts’, imposing an alignment constraint on the prompts produced by Q-Former. However, both Q- and P-Former require a separate PT process for initialization.
2.3 LLM Backbone
Taking LLMs Zhao et al. (2023c); Naveed et al. (2023); Luo et al. (2023) as the core agents, MM-LLMs can inherit some notable properties like zero-shot generalization, few-shot ICL, Chain-of-Thought (CoT), and instruction following. The LLM Backbone processes representations from various modalities, engaging in semantic understanding, reasoning, and decision-making regarding the inputs. It produces (1) direct textual outputs , and (2) signal tokens from other modalities (if any). These signal tokens act as instructions to guide the generator on whether to produce MM contents and, if affirmative, specifying the content to produce:
where the aligned representations of other modalities can be considered as soft Prompt-tuning for the LLM Backbone. Moreover, some research works have introduced Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix-tuning Li and Liang (2021), Adapter Houlsby et al. (2019), and LoRA Hu et al. (2021). In these cases, the number of additional trainable parameters is exceptionally minimal, even less than 0.1% of the total LLM parameter count. We provide an introduction to mainstream PEFT methods in Appendix B.
The commonly used LLMs in MM-LLMs incude Flan-T5 Chung et al. (2022), ChatGLM Zeng et al. (2022a), UL2 Tay et al. (2022), Qwen Bai et al. (2023a), Chinchilla Hoffmann et al. (2022), OPT Zhang et al. (2022b), PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), and Vicuna Chiang et al. (2023). We provide a brief introduction to these models in Appendix C.
2.4 Output Projector
The Output Projector maps the signal token representations from the LLM Backbone into features understandable to the following Modality Generator . Given the -text dataset , is first fed into LLM to generate the corresponding , then mapped into . To facilitate alignment of the mapped features , the goal is to minimize the distance between and the conditional text representations of :
The optimization only relies on captioning texts, without utilizing any audio or visual resources , where and is the textual condition encoder in . The Output Projector is implemented by a Tiny Transformer or MLP.
2.5 Modality Generator
The Modality Generator is tasked with producing outputs in distinct modalities. Commonly, existing works use off-the-shelf Latent Diffusion Models (LDMs) Zhao et al. (2022), i.e., Stable Diffusion Rombach et al. (2022) for image synthesis, Zeroscope Cerspense (2023) for video synthesis, and AudioLDM-2 Liu et al. (2023b, c) for audio synthesis. The features mapped by the Output Projector serve as conditional inputs in the denoising process to generate MM content. During training, the ground truth content is first transformed into a latent feature by the pre-trained VAE Kingma and Welling (2013). Then, noise is added to to obtain the noisy latent feature . A pre-trained Unet Ronneberger et al. (2015) is used to compute the conditional LDM loss as follows:
optimize parameters and by minimizing .
3 Training Pipeline
MM-LLMs’ training pipeline can be delineated into two principal stages: MM PT and MM IT.
3.1 MM PT
During the PT stage, typically leveraging the X-Text datasets, Input and Output Projectors are trained to achieve alignment among various modalities by optimizing predefined objectives (PEFT is sometimes applied to the LLM Backbone). For MM understanding models, optimization focuses solely on Equation (2), while for MM generation models, optimization involves Equations (2), (4), and (5). In the latter case, Equation (2) also includes the ground-truth signal token sequence.
The X-Text datasets encompass Image-Text, Video-Text, and Audio-Text, with Image-Text having two types: Image-Text pairs (i.e., <img1><txt1>) and interleaved Image-Text corpus (i.e., <txt1><img1><txt2><txt3><img2><txt4>). The detailed statistics for these X-Text datasets are presented in Table 3 of Appendix F.
3.2 MM IT
MM IT is a methodology that entails the fine-tuning of pre-trained MM-LLMs using a set of instruction-formatted datasets Wei et al. (2021). Through this tuning process, MM-LLMs can generalize to unseen tasks by adhering to new instructions, thereby enhancing zero-shot performance. This straightforward yet impactful concept has catalyzed the success of subsequent endeavors in the field of NLP, exemplified by works such as InstructGPT Ouyang et al. (2022), OPT-IML Iyer et al. (2022), and InstructBLIP Dai et al. (2023).
MM IT comprises Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), aiming to align with human intents or preferences and enhance the interaction capabilities of MM-LLMs. SFT converts part of the PT stage data into an instruction-aware format. Using visual Question-Answer (QA) as an example, various templates may be employed like (1) <Image>{Question} A short answer to the question is; (2) <Image>Examine the image and respond to the following question with a brief answer: {Question}. Answer:; and so on. Next, it fine-tunes the pre-trained MM-LLMs using the same optimization objectives. The SFT dataset can be structured as either single-turn QA or multi-turn dialogues.
After SFT, RLHF involves further fine-tuning of the model, relying on feedback regarding the MM-LLMs’ responses (e.g., Natural Language Feedback (NLF) labeled manually or automatically) Sun et al. (2023). This process employs a reinforcement learning algorithm to effectively integrate the non-differentiable NLF. The model is trained to generate corresponding responses conditioned on the NLF Chen et al. (2023h); Akyürek et al. (2023). The statistics for SFT and RLHF datasets are presented in Table 4 of Appendix F.
The datasets used by existing MM-LLMs in the MM PT and MM IT stages are diverse, but they are all subsets of the datasets in Tables 3 and 4.
| Model | IO | Modality Encoder | Input Projector | LLM Backbone | Output Projector | Modality Generator | #.PT | #.IT |
| Flamingo | I+V+TT | I/V: NFNet-F6 | Cross-attention | Chinchilla-1.4B/7B/70B (Frozen) | – | – | – | – |
| BLIP-2 | I+TT | I: CLIP/Eva-CLIP ViT@224 | Q-Former w/ Linear Projector | Flan-T5/OPT (Frozen) | – | – | 129M | – |
| LLaVA | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/13B (PT: Frozen; IT: PEFT) | – | – | – | – |
| MiniGPT-4 | I+TT | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-13B (PT: Frozen; IT: PEFT) | – | – | – | – |
| mPLUG-Owl | I+TT | I: CLIP ViT-L/14 | Cross-attention | LLaMA-7B(PT: Frozen; IT: PEFT) | – | – | – | – |
| X-LLM | I+V+A+TT | I/V: ViT-G; A: C-Former | Q-Former w/ Linear Projector | ChatGLM-6B (Frozen) | – | – | – | – |
| VideoChat | V+TT | I: ViT-G | Q-Former w/ Linear Projector | Vicuna (Frozen) | – | – | – | – |
| InstructBLIP | I+V+TT | I/V: ViT-G/14@224 | Q-Former w/ Linear Projector | Flan-T5/Vicuna (Frozen) | – | – | 129M | 1.2M |
| PandaGPT | I+TT | I: ImageBind | Linear Projector | Vicuna-13B (PEFT) | – | – | – | – |
| PaLI-X | I+TT | I: ViT | Linear Projector | UL2-32B (PEFT) | – | – | – | – |
| Video-LLaMA | I+V+A+TT | I/V: EVA-CLIP ViT-G/14; A: ImageBind | Q-Former w/ Linear Projector | Vicuna/LLaMA (Frozen) | – | – | – | – |
| Video-ChatGPT | V+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-v1.1 (Initialized with LLaVA, Frozen) | – | – | – | – |
| Shikra | I+TT | I: CLIP ViT-L/14@224 | Linear Projector | Vicuna-7B/13B (PEFT) | – | – | 600K | 5.5M |
| DLP | I+TT | I: CLIP/Eva-CLIP ViT | Q-Former+P-Former w/ Linear Projector | OPT/Flan-T5 (Frozen) | – | – | – | – |
| BuboGPT | I+A+TT | I: CLIP/Eva-CLIP ViT; A: ImageBind | Q-Former w/ Linear Projector | Vicuna (Frozen) | – | – | – | – |
| ChatSpot | I+TT | I: CLIP ViT-L/14 | Linear Projector | Vicuna-7B/LLaMA (PT: Frozen; IT: PEFT) | – | – | – | – |
| Qwen-VL-(Chat) | I+TT | I: ViT@448 initialized from OpenClip’s ViT-bigG | Cross-attention | Qwen-7B (PT: Frozen; IT: PEFT) | – | – | 1.4B | 50M |
| NExT-GPT | I+V+A+TI+V+A+T | I/V/A: ImageBind | Linear Projector | Vicuna-7B (PEFT) | Tiny Transformer | I: Stable Diffusion; V: Zeroscope; A: AudioLDM | – | – |
| MiniGPT-5 | I+TI+T | I: Eva-CLIP ViT-G/14 | Q-Former w/ Linear Projector | Vicuna-7B (PEFT) | Tiny Transformer w/ MLP | I: StableDiffusion-2 | – | – |
| LLaVA-1.5 | I+TT | I: CLIP ViT-L@336 | MLP | Vicuna-v1.5-7B/13B (PT: Frozen; IT: PEFT) | – | – | 0.6M | 0.7M |
| MiniGPT-v2 | I+TT | I: Eva-CLIP ViT@448 | Linear Projector | LLaMA-2-Chat-7B (PEFT) | – | – | – | – |
| CogVLM | I+TT | I: Eva-2-CLIP ViT | MLP | Vicuna-v1.5-7B (PEFT) | – | – | – | – |
| DRESS | I+TT | I:Eva-CLIP ViT-G/14 | Linear Projector | Vicuna-v1.5-13B (PEFT) | – | – | – | – |
| X-InstructBLIP | I+V+A+3D+TT | I/V: Eva-CLIP ViT-G/14; A: BEATs; 3D: ULIP-2 | Q-Former w/ Linear Projector | Vicuna-v1.1-7B/13B (Frozen) | – | – | – | – |
| CoDi-2 | I+V+A+TI+V+A+T | I/V/A: ImageBind | MLP | LLaMA-2-Chat-7B (PT: Frozen; IT: PEFT) | MLP | I: Stable Diffusion-2.1; V: Zeroscope-v2; A: AudioLDM-2 | – | – |
| VILA | I+TT | I: ViT@336 | Linear Projector | LLaMA-2-7B/13B (PEFT) | – | – | 50M | 1M |
4 SOTA MM-LLMs
Based on the previously defined design formulations, we conduct a comprehensive comparison of the architectures and training dataset scales for SOTA MM-LLMs, as illustrated in Table 1. Subsequently, we will provide a concise introduction to the core contributions of these models and summarize their developmental trends.
(1) Flamingo Alayrac et al. (2022) represents a series of Visual Language (VL) Models designed for processing interleaved visual data and text, generating free-form text as the output. (2) BLIP-2 Li et al. (2023c) introduces a more resource-efficient framework, comprising the lightweight Q-Former to bridge modality gaps and the utilization of frozen LLMs. Leveraging LLMs, BLIP-2 can be guided for zero-shot image-to-text generation using natural language prompts. (3) LLaVA Liu et al. (2023e) pioneers the transfer of IT techniques to the MM domain. Addressing data scarcity, LLaVA introduces a novel open-source MM instruction-following dataset created using ChatGPT/GPT-4, alongside the MM instruction-following benchmark, LLaVA-Bench. (4) MiniGPT-4 Zhu et al. (2023a) proposes a streamlined approach where training only one linear layer aligns the pre-trained vision encoder with the LLM. This efficient method enables the replication of the exhibited capabilities of GPT-4. (5) mPLUG-Owl Ye et al. (2023) presents a novel modularized training framework for MM-LLMs, incorporating the visual context. To assess different models’ performance in MM tasks, the framework includes an instructional evaluation dataset called OwlEval. (6) X-LLM Chen et al. (2023b) is expanded to various modalities, including audio, and demonstrates strong scalability. Leveraging the language transferability of the Q-Former, X-LLM is successfully applied in the context of Sino-Tibetan Chinese. (7) VideoChat Li et al. (2023d) pioneers an efficient chat-centric MM-LLM for video understanding dialogue, setting standards for future research in this domain and offering protocols for both academia and industry. (8) InstructBLIP Dai et al. (2023) is trained based on the pre-trained BLIP-2 model, updating only the Q-Former during MM IT. By introducing instruction-aware visual feature extraction and corresponding instructions, the model enables the extraction of flexible and diverse features. (9) PandaGPT Su et al. (2023) is a pioneering general-purpose model with the capability to comprehend and act upon instructions across different modalities: text, image/video, audio, thermal, depth, and inertial measurement units. (10) PaLI-X Chen et al. (2023g) is trained using mixed VL objectives and unimodal objectives, including prefix completion and masked-token completion. This approach proves effective for both downstream task results and achieving the Pareto frontier in the fine-tuning setting. (11) Video-LLaMA Zhang et al. (2023e) introduces a multi-branch cross-modal PT framework, enabling LLMs to simultaneously process the vision and audio content of a given video while engaging in conversations with humans. This framework aligns vision with language as well as audio with language. (12) Video-ChatGPT Maaz et al. (2023) is a model specifically designed for video conversations, capable of generating discussions about videos by integrating spatiotemporal vision representations. (13) Shikra Chen et al. (2023d) introduces a simple and unified pre-trained MM-LLM tailored for Referential Dialogue, a task involving discussions about regions and objects in images. This model demonstrates commendable generalization ability, effectively handling unseen settings. (14) DLP Jian et al. (2023) proposes the P-Former to predict the ideal prompt, trained on a dataset of single-modal sentences. This showcases the feasibility of single-modal training to enhance MM learning. (15) BuboGPT Zhao et al. (2023d) is a model constructed by learning a shared semantic space for a comprehensive understanding of MM content. It explores fine-grained relationships among different modalities such as image, text, and audio. (16) ChatSpot Zhao et al. (2023b) introduces a simple yet potent method for finely adjusting precise referring instructions for MM-LLM, facilitating fine-grained interactions. The incorporation of precise referring instructions, consisting of image- and region-level instructions, enhances the integration of multi-grained VL task descriptions. (17) Qwen-VL Bai et al. (2023b) is a multi-lingual MM-LLM that supports both English and Chinese. Qwen-VL also allows the input of multiple images during the training phase, improving its ability to understand the vision context. (18) NExT-GPT Wu et al. (2023d) is an end-to-end, general-purpose any-to-any MM-LLM that supports the free input and output of image, video, audio, and text. It employs a lightweight alignment strategy, utilizing LLM-centric alignment in the encoding phase and instruction-following alignment in the decoding phase. (19) MiniGPT-5 Zheng et al. (2023b) is an MM-LLM integrated with inversion to generative vokens and integration with Stable Diffusion. It excels in performing interleaved VL outputs for MM generation. The inclusion of classifier-free guidance during the training phase enhances the quality of generation.
| Model | LLM Backbone | OKVQA | IconVQA | VQA | GQA | VizWiz | SQA | VQA | POPE | MME | MME | MMB | MMB | SEED | LLaVA | MM-Vet | QBench | HM | VSR |
| Flamingo | Chinchilla-7B | 44.7 | – | – | – | 28.8 | – | – | – | – | – | – | – | – | – | – | – | 57.0 | 31.8 |
| BLIP-2 | Flan-T5 | 45.9 | 40.6 | 65.0 | 44.7 | 19.6 | 61.0 | 42.5 | 85.3 | 1293.8 | 290.0 | – | – | 46.4 | 38.1 | 22.4 | – | 53.7 | 50.9 |
| LLaVA | Vicuna-13B | 54.4 | 43.0 | – | 41.3 | – | – | 38.9 | – | – | – | – | – | – | – | – | – | – | 51.2 |
| MiniGPT-4 | Vicuna-13B | 37.5 | 37.6 | – | 30.8 | – | – | 19.4 | – | – | – | – | – | – | – | – | – | – | 41.6 |
| InstructBLIP | Vicuna-7B | – | – | – | 49.2 | 34.5 | 60.5 | 50.1 | – | – | – | 36.0 | 23.7 | 53.4 | 60.9 | 26.2 | 56.7 | – | – |
| InstructBLIP | Vicuna-13B | – | 44.8 | – | 49.5 | 33.4 | 63.1 | 50.7 | 78.9 | 1212.8 | 291.8 | – | – | – | 58.2 | 25.6 | – | 57.5 | 52.1 |
| Shikra | Vicuna-13B | 47.2 | – | 77.4 | – | – | – | – | – | – | – | 58.8 | – | – | – | – | 54.7 | – | – |
| IDEFICS-9B | LLaMA-7B | – | – | 50.9 | 38.4 | 35.5 | – | 25.9 | – | – | – | 48.2 | 25.2 | – | – | – | – | – | – |
| IDEFICS-80B | LLaMA-65B | – | – | 60.0 | 45.2 | 36.0 | – | 30.9 | – | – | – | 54.5 | 38.1 | – | – | – | – | – | – |
| Qwen-VL | Qwen-7B | – | – | 78.8 | 59.3 | 35.2 | 67.1 | 63.8 | – | – | – | 38.2 | 7.4 | 56.3 | – | – | 59.4 | – | – |
| Qwen-VL-Chat | Qwen-7B | – | – | 78.2 | 57.5 | 38.9 | 68.2 | 61.5 | – | 1487.5 | 360.7 | 60.6 | 56.7 | 58.2 | – | – | – | – | – |
| LLaVA-1.5 | Vicuna-1.5-7B | – | – | 78.5 | 62.0 | 50.0 | 66.8 | 58.2 | 85.9 | 1510.7 | 316.1 | 64.3 | 58.3 | 58.6 | 63.4 | 30.5 | 58.7 | – | – |
| +ShareGPT4V | Vicuna-1.5-7B | – | – | 80.6 | – | 57.2 | 68.4 | – | – | 1567.4 | 376.4 | 68.8 | 62.2 | 69.7 | 72.6 | 37.6 | 63.4 | – | – |
| LLaVA-1.5 | Vicuna-1.5-13B | – | – | 80.0 | 63.3 | 53.6 | 71.6 | 61.3 | 85.9 | 1531.3 | 295.4 | 67.7 | 63.6 | 61.6 | 70.7 | 35.4 | 62.1 | – | – |
| MiniGPT-v2 | LLaMA-2-Chat-7B | 56.9 | 47.7 | – | 60.3 | 30.3 | – | 51.9 | – | – | – | – | – | – | – | – | – | 58.2 | 60.6 |
| MiniGPT-v2-Chat | LLaMA-2-Chat-7B | 55.9 | 49.4 | – | 58.8 | 42.4 | – | 52.3 | – | – | – | – | – | – | – | – | – | 59.5 | 63.3 |
| VILA-7B | LLaMA-2-7B | – | – | 79.9 | 62.3 | 57.8 | 68.2 | 64.4 | 85.5 | 1533.0 | – | 68.9 | 61.7 | 61.1 | 69.7 | 34.9 | – | – | – |
| VILA-13B | LLaMA-2-13B | – | – | 80.8 | 63.3 | 60.6 | 73.7 | 66.6 | 84.2 | 1570.1 | – | 70.3 | 64.3 | 62.8 | 73.0 | 38.8 | – | – | – |
| +ShareGPT4V | LLaMA-2-13B | – | – | 80.6 | 63.2 | 62.4 | 73.1 | 65.3 | 84.8 | 1556.5 | – | 70.8 | 65.4 | 61.4 | 78.4 | 45.7 | – | – | – |
For introduction regarding the remaining seven MM-LLMs, please refer to Appendix D, which includes (20) LLaVA-1.5 Liu et al. (2023d), (21) MiniGPT-v2 Chen et al. (2023c), (22) CogVLM Wang et al. (2023), (23) DRESS Chen et al. (2023h), (24) X-InstructBLIP Panagopoulou et al. (2023), (25) CoDi-2 Tang et al. (2023a), and (26) VILA Lin et al. (2023).
Trends in Existing MM-LLMs:
(1) Progressing from a dedicated emphasis on MM understanding to the generation of specific modalities and further evolving into any-to-any modality conversion (e.g., MiniGPT-4 MiniGPT-5 NExT-GPT); (2) Advancing from MM PT to SFT and then to RLHF, the training pipeline undergoes continuous refinement, striving to better align with human intent and enhance the model’s conversational interaction capabilities (e.g., BLIP-2 InstructBLIP DRESS); (3) Embracing Diversified Modal Extensions (e.g., BLIP-2 X-LLM and InstructBLIP X-InstructBLIP); (4) Incorporating a Higher-Quality Training Dataset (e.g., LLaVA LLaVA-1.5); (5) Adopting a More Efficient Model Architecture, transitioning from complex Q- and P-Former input projector modules in BLIP-2 and DLP to a simpler yet effective linear projector in VILA.
5 Benckmarks and Performance
To offer a comprehensive performance comparison, we have compiled a table featuring major MM-LLMs across VL benchmarks gathered from various papers Li et al. (2023c); Chen et al. (2023c, e); Lin et al. (2023), as shown in Table 2. The information of these benchmarks can be found in Appendix E. Next, we will extract essential training recipes that boost the effectiveness of MM-LLMs, drawing insights from SOTA models.
Training Recipes
Firstly, higher image resolution can incorporate more visual details for the model, benefiting tasks that require fine-grained details. For example, LLaVA-1.5 and VILA employ a resolution of , while Qwen-VL and MiniGPT-v2 utilize . However, higher resolutions lead to longer token sequences, incurring additional training and inference costs. MiniGPT-v2 addresses this by concatenating adjacent visual tokens in the embedding space to reduce length. Recently, Monkey Li et al. (2023h) proposed a solution to enhance the resolution of input images without retraining a high-resolution visual encoder, utilizing only a low-resolution visual encoder, supporting resolutions up to . To enhance the understanding of rich-text images, tables, and document content, DocPedia Feng et al. (2023) introduced a method to increase the visual encoder resolution to , overcoming the limitations of poorly performing low resolutions in open-sourced ViT. Secondly, the incorporation of high-quality SFT data can significantly improve performance in specific tasks, as evidenced by the addition of ShareGPT4V data to LLaVA-1.5 and VILA-13B, as shown in Table 2. Moreover, VILA reveals several key findings: (1) Performing PEFT on the LLM Backbone promotes deep embedding alignment, crucial for ICL; (2) Interleaved Image-Text data proves beneficial, whereas Image-Text pairs alone are sub-optimal; (3) Re-blending text-only instruction data (e.g., Unnatural instruction Honovich et al. (2022)) with image-text data during SFT not only addresses the degradation of text-only tasks but also enhances VL task accuracy.
6 Future Directions
In this section, we explore promising future directions for MM-LLMs across the following aspects:
More Powerful Models
We can enhance the MM-LLMs’ strength from the following four key avenues: (1) Expanding Modalities: Current MM-LLMs typically support the following modalities: image, video, audio, 3D, and text. However, the real world involves a broader range of modalities. Extending MM-LLMs to accommodate additional modalities (e.g., web pages, heat maps, and figures&tables) will increase the model’s versatility, making it more universally applicable; (2) Diversifying LLMs: Incorporating various types and sizes of LLMs provides practitioners with the flexibility to select the most appropriate one based on their specific requirements; (3) Improving MM IT Dataset Quality: Current MM IT dataset have ample room for improvement and expansion. Diversifying the range of instructions can enhance the effectiveness of MM-LLMs in understanding and executing user commands. (4) Strengthening MM Generation Capabilities: Most current MM-LLMs are predominantly oriented towards MM understanding. Although some models have incorporated MM generation capabilities, the quality of generated responses may be constrained by the capacities of the LDMs. Exploring the integration of retrieval-based approaches Asai et al. (2023) holds significant promise in complementing the generative process, potentially enhancing the overall performance of the model.
More Challenging Benchmarks
Existing benchmarks might not adequately challenge the capabilities of MM-LLMs, given that many datasets have previously appeared to varying degrees in the PT or IT sets. This implies that the models may have learned these tasks during training. Moreover, current benchmarks predominantly concentrate on the VL sub-field. Thus, it is crucial for the development of MM-LLMs to construct a more challenging, larger-scale benchmark that includes more modalities and uses a unified evaluation standard. Concurrently, benchmarks can be tailored to assess the MM-LLMs’ proficiency in practical applications. For instance, the introduction of GOAT-Bench Lin et al. (2024) aims to evaluate various MM-LLMs’ capacity to discern and respond to nuanced aspects of social abuse presented in memes.
Mobile/Lightweight Deployment
To deploy MM-LLMs on resource-constrained platforms and achieve optimal performance meanwhile, such as low-power mobile and IoT devices, lightweight implementations are of paramount importance. A notable advancement in this realm is MobileVLM Chu et al. (2023a). This approach strategically downscales LLaMA, allowing for seamless off-the-shelf deployment. MobileVLM further introduces a Lightweight Downsample Projector, consisting of fewer than million parameters, contributing to improved computational speed. Nevertheless, this avenue necessitates additional exploration for further advancements in development.
Embodied Intelligence
The embodied intelligence aims to replicate human-like perception and interaction with the surroundings by effectively understanding the environment, recognizing pertinent objects, assessing their spatial relationships, and devising a comprehensive task plan Firoozi et al. (2023). Embodied AI tasks, such as embodied planning, embodied visual question answering, and embodied control, equips robots to autonomously implement extended plans by leveraging real-time observations. Some typical work in this area is PaLM-E Driess et al. (2023) and EmbodiedGPT Mu et al. (2023). PaLM-E introduces a multi-embodiment agent through the training of a MM-LLM. Beyond functioning solely as an embodied decision maker, PaLM-E also demonstrates proficiency in handling general VL tasks. EmbodiedGPT introduces an economically efficient method characterized through a CoT approach, enhancing the capability of embodied agents to engage with the real world and establishing a closed loop that connects high-level planning with low-level control. While MM-LLM-based Embodied Intelligence has made advancements in integrating with robots, further exploration is needed to enhance the autonomy of robots.
Continual IT
In practical applications, MM-LLMs are expected to adapt to new MM tasks for supporting additional functionalities. Nevertheless, current MM-LLMs remain static and are unable to adjust to continuously emerging requirements. Therefore, an approach is needed to make the model flexible enough to efficiently and continually leverage emerging data, while avoiding the substantial cost of retraining MM-LLMs. This aligns with the principles of continual learning, where models are designed to incrementaly learn new tasks similar to human learning. Continual IT aims to continuously fine-tune MM-LLMs for new MM tasks while maintaining superior performance on tasks learned during the original MM IT stage. It introduces two primary challenges: (1) catastrophic forgetting, where models forget previous knowledge when learning new tasks Robins (1995); McCloskey and Cohen (1989); Goodfellow et al. (2013); Zhang et al. (2023d, c, b); Zheng et al. (2023a), and (2) negative forward transfer, indicating that the performance of unseen tasks is declined when learning new ones Zheng et al. (2024); Dong et al. (2023b, a). Recently, He et al. established a benchmark to facilitate the development of continual IT for MM-LLMs. Despite these advancements, there is still a significant opportunity and room for improvement in developing better methods to address the challenges of catastrophic forgetting and negative forward transfer.
7 Conclusion
In this paper, we have presented a comprehensive survey of MM-LLMs with a focus on recent advancements. Initially, we categorize the model architecture into five components, providing a detailed overview of general design formulations and training pipelines. Subsequently, we introduce various SOTA MM-LLMs, each distinguished by its specific formulations. Our survey also sheds light on their capabilities across diverse MM benchmarks and envisions future developments in this rapidly evolving field. We hope this survey can provide insights for researchers, contributing to the ongoing advancements in the MM-LLMs domain.
Limitations
In this paper, we embark on a comprehensive exploration of the current MM-LLMs landscape, presenting a synthesis from diverse perspectives enriched by our insights. Acknowledging the dynamic nature of this field, it is plausible that certain aspects may have eluded our scrutiny, and recent advances might not be entirely encapsulated. To tackle this inherent challenge, we’ve established a dedicated website for real-time tracking, using crowdsourcing to capture the latest advancements. Our goal is for this platform to evolve into a continuous source of contributions propelling ongoing development in the field. Given the constraints of page limits, we are unable to delve into all technical details and have provided concise overviews of the core contributions of mainstream MM-LLMs. Looking ahead, we commit to vigilant monitoring and continual enhancement of relevant details on our website, incorporating fresh insights as they emerge.
Appendix A Related Surveys
Prior to the emergence of LLMs, several surveys on traditional MM PT have been conducted Ruan and Jin (2022); Du et al. (2022a); Long et al. (2022); Chen et al. (2023a). Most of these models entail a substantial computational cost during the PT phase, attributable to end-to-end training using large-scale models and datasets. As a consequence of not incorporating LLMs, these models suffer from deficiencies in instruction following, ICL, CoT, and interactive capabilities. Moreover, the training pipeline solely encompasses the PT phase without the inclusion of an IT stage.
In recent times, several surveys have emerged on MM-LLMs. Yin et al. and Wu et al. exclusively delve into early VL understanding models. Huang et al. place a primary emphasis on visual IT, while Song et al. focus on modal alignment methods. Lastly, Cui et al. provide a comprehensive review of the applications of MM-LLMs within the realm of autonomous driving.
Compared with their works, the main distinctions are outlined as follows:
We have comprehensively covered nearly all MM-LLMs over the past year, including not only understanding models but also generative models. Our coverage extends beyond VL modalities to encompass various modes such as audio and 3D;
To offer readers a comprehensive understanding of MM-LLMs, we have introduced a general model architecture that incorporates any-to-any modality transformations, offering a detailed overview of the functional roles and implementation choices for each component;
We have summarized the developmental trends of existing MM-LLMs and provided some training recipes that can enhance effectiveness;
We have established an open-source website for MM-LLMs researchers, supporting crowdsourced updates and aiming to facilitate collaboration in the MM-LLMs field. We anticipate that this survey will illuminate future research in the MM-LLMs domain.
Appendix B Mainstream PEFT Methods
PEFT entails maintaining the pre-trained LLM in a frozen state while adjusting a small number of additional trainable parameters. In the following section, we revisit several representative PEFT methods, where and represent the input and output of the original module, and signifies the output of this module when attached with PEFT.
Prefix-tuning
Adapter
Houlsby et al. (2019); He et al. (2021); Rebuffi et al. (2017); Zhang et al. (2020) is typically a residual block consisting of a down-projection matrix , a nonlinear activation function , and an up-projection matrix . It can be inserted into any layer of the pre-trained LLM, formulated as follows:
LoRA
Hu et al. (2021) is the most commonly used PEFT method. It assumes that the change in parameters occurs within a low-rank space. Given a pre-trained matrix , LoRA learns an incremental update and decomposes into a matrix multiplication between two low-rank matrices and , where min. LoRA follows the forward process as outlined below:
QLoRA Dettmers et al. (2023) is a quantized LoRA. The underlying principle of QLoRA includes the quantization of pre-trained weights to bits, followed by the execution of PEFT using LoRA.
Appendix C Commonly Used LLMs
The commonly used LLM Backbones in existing MM-LLMs research are as follows:
ChatGLM222https://github.com/THUDM/ChatGLM-6B is a Chinese-English bilingual dialogue model, optimized by an auto-regressive mask infilling objective. It is based on the GLM Du et al. (2022b); Zeng et al. (2022a) architecture, optimized for Chinese question answering and dialogues.
UL2 Tay et al. (2022) is an encoder-decoder model trained utilizing a mixture of denoisers objectives, surpassing T5 on numerous benchmarks.
Qwen Bai et al. (2023a) is trained on large-scale and diverse datasets, with a primary focus on Chinese and English. It employs SFT and RLHF techniques for alignment, resulting in dialogue models like Qwen-Chat.
Chinchilla Hoffmann et al. (2022) is a causal decoder, trained on extensive text data. It posits that model size should double for every doubling of training tokens.
PaLM Chowdhery et al. (2023) is a causal decoder structure with parallel attention and feed-forward layers, enabling training speeds up to times faster. Notable changes contain RoPE embeddings, SwiGLU activation, multi-query attention, and etc.
LLaMA Touvron et al. (2023a) comprises decoder-only models with efficient causal attention.
LLaMA-2 Touvron et al. (2023b) focuses on fine-tuning a superior and safer LLaMA-2-Chat model for conversation generation, incorporating 40% more training data with grouped-query attention and a larger context length.
Vicuna Chiang et al. (2023) is a model built on top of LLaMA, utilizing user dialogue data obtained from ShareGPT.com and trained by SFT.
Appendix D SOTA MM-LLMs (continued)
(20) LLaVA-1.5 Liu et al. (2023d) reports simple modifications to the LLaVA framework, including applying an MLP projection and introducing VQA data tailored for academic tasks, along with simple response formatting prompts. These adjustments result in enhanced capabilities for MM understanding.
(21) MiniGPT-v2 Chen et al. (2023c) is an MM-LLM designed as a unified interface for diverse VL multi-task learning. To create a single model proficient in handling multiple VL tasks, identifiers are incorporated for each task during both training and inference. This facilitates clear task distinction, ultimately enhancing learning efficiency.
(22) CogVLM Wang et al. (2023) is an open-source MM-LLM that bridges the gap between modalities via a trainable visual expert module within the attention and feedforward layers. This allows for a deep fusion of MM features without compromising performance on NLP downstream tasks.
(23) DRESS Chen et al. (2023h) introduces a method using natural language feedback to enhance alignment with human preferences. DRESS extends the conditional reinforcement learning algorithm to integrate non-differentiable natural language feedback, training the model to generate appropriate responses based on feedback.
(24) X-InstructBLIP Panagopoulou et al. (2023) introduces a cross-modal framework with instruction-aware representations, scalable enough to empower LLMs to handle diverse tasks across multiple modalities, including image/video, audio, and 3D. Notably, it achieves this without the need for modality-specific PT.
(25) CoDi-2 Tang et al. (2023a) is a MM generation model excelling in modality-interleaved instruction following, in-context generation, and user-model interaction by multi-turn conversations. It enhances CoDi Tang et al. (2023b) to process intricate modality-interleaved inputs and instructions, generating latent features autoregressively.
(26) VILA Lin et al. (2023) outperforms in vision tasks and shows remarkable reasoning ability while maintaining text-only capabilities. It achieves this by harnessing the full capabilities of LLM learning, using the interleaved attributes of image-text pairs, and implementing meticulous text data re-blending.
| Dataset Name | X Modality | #.X | #.T | #.X-T |
| ALIGN Jia et al. (2021) | Image | 1.8B | 1.8B | 1.8B |
| LTIP Alayrac et al. (2022) | Image | 312M | 312M | 312M |
| MS-COCO Lin et al. (2014) | Image | 124K | 620K | 620K |
| Visual Genome Krishna et al. (2017) | Image | 108K | 4.5M | 4.5M |
| CC3M Sharma et al. (2018) | Image | 3.3M | 3.3M | 3.3M |
| CC12M Changpinyo et al. (2021) | Image | 12.4M | 12.4M | 12.4M |
| SBU Ordonez et al. (2011) | Image | 1M | 1M | 1M |
| LAION-5B Schuhmann et al. (2022) | Image | 5.9B | 5.9B | 5.9B |
| LAION-400M Schuhmann et al. (2021) | Image | 400M | 400M | 400M |
| LAION-en Schuhmann et al. (2022) | Image | 2.3B | 2.3B | 2.3B |
| LAION-zh Schuhmann et al. (2022) | Image | 142M | 142M | 142M |
| LAION-COCO Schuhmann et al. (2022b) | Image | 600M | 600M | 600M |
| Flickr30k Young et al. (2014) | Image | 31K | 158K | 158K |
| AI Challenger Captions Wu et al. (2017) | Image | 300K | 1.5M | 1.5M |
| COYO Byeon et al. (2022) | Image | 747M | 747M | 747M |
| Wukong Gu et al. (2022) | Image | 101M | 101M | 101M |
| COCO Caption Chen et al. (2015) | Image | 164K | 1M | 1M |
| WebLI Chen et al. (2022b) | Image | 10B | 12B | 12B |
| Episodic WebLI Chen et al. (2023g) | Image | 400M | 400M | 400M |
| CC595k Liu et al. (2023e) | Image | 595K | 595K | 595K |
| RefCOCO Kazemzadeh et al. (2014) | Image | 20K | 142K | 142K |
| RefCOCO+ Yu et al. (2016) | Image | 20K | 142K | 142K |
| Visual-7W Zhu et al. (2016) | Image | 47.3K | 328K | 328K |
| OCR-VQA Mishra et al. (2019) | Image | 207K | 1M | 1M |
| ST-VQA Biten et al. (2022) | Image | 23K | 32K | 32K |
| DocVQA Mathew et al. (2021) | Image | 12K | 50K | 50K |
| TextVQA Singh et al. (2019) | Image | 28.4K | 45.3K | 45.3K |
| DataComp Gadre et al. (2023) | Image | 1.4B | 1.4B | 1.4B |
| GQA Hudson and Manning (2019) | Image | 113K | 22M | 22M |
| VGQA Krishna et al. (2017) | Image | 108K | 1.7M | 1.7M |
| VQA Goyal et al. (2017) | Image | 265K | 1.4M | 1.4M |
| DVQA Kafle et al. (2018) | Image | 300K | 3.5M | 3.5M |
| OK-VQA Schwenk et al. (2022) | Image | 14K | 14K | 14K |
| A-OKVQA Schwenk et al. (2022) | Image | 23.7K | 24.9K | 24.9K |
| Text Captions Sidorov et al. (2020) | Image | 28K | 145K | 145K |
| M3W (Interleaved) Alayrac et al. (2022) | Image | 185M | 182GB | 43.3M (Instances) |
| MMC4 (Interleaved) Zhu et al. (2023b) | Image | 571M | 43B | 101.2M (Instances) |
| MSRVTT Xu et al. (2016) | Video | 10K | 200K | 200K |
| WebVid Bain et al. (2021) | Video | 10M | 10M | 10M |
| VTP Alayrac et al. (2022) | Video | 27M | 27M | 27M |
| AISHELL-2 Chen et al. (2023b) | Audio | – | – | 128K |
| AISHELL-2 Chen et al. (2023b) | Audio | – | – | 1M |
| WaveCaps Mei et al. (2023) | Audio | 403K | 403K | 403K |
| VSDial-CN Chen et al. (2023b) | Image, Audio | 120K (Image), 1.2M(Audio) | 120K | 1.2M |
| Dataset Name | Type | IO | Source | Method | Multi-Turn | #.I/V/A | #.Dialog Turn | #.Instance |
| MiniGPT-4’s IT Zhu et al. (2023a) | SFT | I+TT | CC3M, CC12M | Auto. | ✗ | 134M/–/– | 1 | 5K |
| StableLLaVA Li et al. (2023f) | SFT | I+TT | SD Rombach et al. (2022) | Auto.+Manu. | ✗ | 126K/–/– | 1 | 126K |
| LLaVA’s IT Zhang et al. (2023f) | SFT | I+TT | MS-COCO | Auto. | ✔ | 81K/–/– | 2.29 | 150K |
| SVIT Zhao et al. (2023a) | SFT | I+TT | MS-COCO, Visual Genome | Auto. | ✔ | 108K/–/– | 5 | 3.2M |
| LLaVAR Zhang et al. (2023f) | SFT | I+TT | MS-COCO, CC3M, LAION | LLaVA+Auto. | ✔ | 20K/–/– | 2.27 | 174K |
| ShareGPT4V Chen et al. (2023e) | SFT | I+TT | LCS, COCO, SAM, TextCaps, WikiArt | Auto.+Manu. | ✗ | 100K/–/– | – | – |
| DRESS’s IT Chen et al. (2023h) | SFT | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 193K/–/– | 4 | – |
| VideoChat’s IT Li et al. (2023d) | SFT | V+TT | WebVid | Auto. | ✔ | –/8K/– | 1.82 | 11K |
| Video-ChatGPT’s IT Maaz et al. (2023) | SFT | V+TT | ActivityNet Caba Heilbron et al. (2015) | Inherit | ✔ | –/100K/– | 1 | 100K |
| Video-LLaMA’s IT Zhang et al. (2023e) | SFT | I/V+TT | MiniGPT-4, LLaVA, and VideoChat’s IT | Auto. | ✔ | 81K/8K/– | 2.22 | 171K |
| InstructBLIP’s IT Dai et al. (2023) | SFT | I/V+TT | Multiple (InstructBLIP’s Figure 2) | Auto. | ✗ | – | – | 1.6M |
| X-InstructBLIP’s IT Panagopoulou et al. (2023) | SFT | I/V/A/3D+TT | Multiple (X-InstructBLIP’s Figure 4) | Auto. | ✗ | – | – | 1.8M |
| MIMIC-IT Li et al. (2023a) | SFT | I/V+TT | Multiple | Auto. | ✗ | 8.1M/502K/– | 1 | 2.8M |
| PandaGPT’s IT Su et al. (2023) | SFT | I+TT | MiniGPT-4 and LLaVA’s IT | Inherit | ✔ | 81K/–/– | 2.29 | 160K |
| MGVLID Zhao et al. (2023b) | SFT | I+B+TT | Multiple | Auto.+Manu. | ✗ | 108K/–/– | – | 108K |
| MIT Li et al. (2023e) | SFT | I/V/B+TT | Multiple | Auto.+Manu. | ✗ | –/–/– | 1 | 2.4M |
| LAMM Yin et al. (2023b) | SFT | I+3D+TT | Multiple | Auto.+Manu. | ✔ | 91K/–/– | 3.27 | 196K |
| BuboGPT’s IT Zhao et al. (2023d) | SFT | (I+A)/A+TT | Clotho, VGGSS | Auto. | ✗ | 5K/–/9K | – | 9K |
| mPLUG-DocOwl’s IT Ye et al. (2023) | SFT | I/Tab/Web+TT | Multiple | Inherit | ✗ | – | – | – |
| T2M Wu et al. (2023d) | SFT | TI/V/A+T | WebVid, CC3M, AudioCap | Auto. | ✗ | 4.9K/4.9K/4.9K | 1 | 14.7K |
| MosIT Wu et al. (2023d) | SFT | I+V+A+TI+V+A+T | Youtube, Google, Flickr30k, Midjourney, etc. | Auto.+Manu. | ✔ | 4K/4K/4K | 4.8 | 5K |
| DRESS’s IT Chen et al. (2023h) | RLHF | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 33K/–/– | 4 | – |
Appendix E VL Benchmarks
The VL benchmarks presented in Table 2 include OKVQA Schwenk et al. (2022), IconVQA Lu et al. (2021), VQA Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQA: ScienceQA-IMG Lu et al. (2022), VQA: TextVQA Singh et al. (2019), POPE Li et al. (2023g), MME: MME Perception Fu et al. (2023), MME: MME Cognition Fu et al. (2023), MMB: MMBenchmark Liu et al. (2023f), MMB: MMBench-Chinese Liu et al. (2023f), SEED: SEED-Bench (Image) Li et al. (2023b), LLaVA: LLaVA-Bench (In-the-Wild) Liu et al. (2023a), MM-Vet Yu et al. (2023), QBench Wu et al. (2023b), HM: HatefulMemes Kiela et al. (2020), and VSR Liu et al. (2023a).